OpenStudy (anonymous):

Solve for x: x^2 + 24x + 90 = 0

OpenStudy (anonymous):

Do you know the quadratic equation?

OpenStudy (anonymous):

(x+12)^2 -144 +90 = 0 (x+12)^2 -54 = 0 (x+12)^2 = 54 x+12 = ±3√6 x = -12 ±3√6